OpenStudy (anonymous):

In the equation x^2 - 8x - 48 = 0, what are the solutions for x?

myininaya (myininaya):

48=6(8)=3(16)=12(4) -48=-12(4) and -8=-12+4 so x^2-8x-48=(x-12)(x+4)=0 implies x=12 and x=-4